OpenStudy (anonymous):

Given f(x) = 2x + 3 and g(x) = 5x – 2, (g ○ f)(x) =

OpenStudy (jhannybean):

g(f(x)) plug in the function f(x) for all values of x in g(x)

OpenStudy (jhannybean):

f(x) = 2x+3 g(x) = 5x-2 g(f(x)) = 5(2x+3) -2 Can you solve this?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

is it 10x+13

OpenStudy (jhannybean):

yes :D

OpenStudy (jhannybean):

good job.

OpenStudy (anonymous):

its right

OpenStudy (anonymous):

???

OpenStudy (jhannybean):

Yep.